OpenStudy (anonymous):

x^2-4x+3=0 solve each equation by completing the square

OpenStudy (anonymous):

You'll need to move the three from the left hand side to the right hand side so that your equation looks like this: \[x^2-4x=-3\] You'll then need to find the perfect square for the left hand side. After finding the third term that completes the square for the left hand side. Set you equation equal to the right and solve by adding and subtracting the remaining term on the left. For example: Complete the square: \[x^2+6x-7=0\] Set the equation to find the perfect square: \[x^2+6x=7\] Find the perfect square and add the term to both sides: \[x^2+6x+9=7+9\] Simplify and set the perfect square factor: \[(x+3)^2=16\] Set to find x by square rooting both sides (remember to plus and minus on the right side): \[x+3=\pm4\] Isolate x. \[x=-3\pm4\] So then: \[x=-3+4 =-1\] \[x=-3-4=-7\]